OpenStudy (kainui):

A spherical raindrop evaporates at a rate proportional to its surface area. Write a differential equation for the volume of the raindrop as a function of time. Halp!

OpenStudy (anonymous):

V = 4/3*pi*3^3 r = (3/4piV)^1/3 dV/dt = finish it off someone

OpenStudy (anonymous):

The surface area is the derivative of volume with respect to the radius, when we talk about spheres.

OpenStudy (anonymous):

dV/dt = -kV^(2/3)

OpenStudy (anonymous):

You know that after the time is passing by(delta(t)) the raindrop will lose volume(delta(V)), which is proportional to the surface of the raindrop 4(pi)(r^2) then: \[\Delta(V)=-a(4\pi r^2)*\Delta(t)\] with a>0. And the minus is there beacause the volume is decreasing.. As long as the volume of the sphere is \[V=\frac{ 4 }{ 3 }\pi r^3\] results that: \[r=(\frac{ 3 }{ 4\pi }V)^{\frac{ 1 }{ 3 }}\] Now we're going into differential: \[\frac{ dV }{ dt }=-kV ^{\frac{ 2 }{ 3 }}\] Where k is a constant >0

OpenStudy (anonymous):

^ yeah that's what I'm saying!

OpenStudy (anonymous):

\[ \frac{dV}{dt} = -k\frac{dV}{dr} \]

OpenStudy (kainui):

The way I solved this I ended up with \[V(t)=\frac{4}{3} \pi (-kt+r_o)^3\] That seem right to you guys?

OpenStudy (kainui):

I am pretty sure I need to take a nap, I'm being ridiculous looking into my old diffeq book unable to solve this laughable problem. Night!

OpenStudy (anonymous):

Well based on what they have, you can so separation of variables: \[ \int V^{3/2}\frac{dV}{dt} dt= \int -k\;dt \]

OpenStudy (anonymous):

\[ \frac 25V^{5/3} = -kt+C \]